OpenStudy (anonymous):

3 6 over 7 obelus divided by 9????

OpenStudy (anonymous):

i wish, unfortunately my professor feels math isnt hard enough without these ridiculous terms.

OpenStudy (jack1):

obelus symbol is: \[.\div \]

OpenStudy (jack1):

so ur equation is: \[\frac{ 27 }{ 7 } \div9\]

OpenStudy (anonymous):

so how could i possibly divide that? i know i need to turn 9 into a fraction and then flip it into a reciprocal

OpenStudy (jack1):

divided by 9 = times by 1/9

OpenStudy (anonymous):

2 over 7?

OpenStudy (jack1):

so same as before, multiplying fractions: times the top terms together over times the bottom terms together =27x1 / 9x7 =?

OpenStudy (anonymous):

27 over 63

OpenStudy (jack1):

27/63 =3/7

OpenStudy (jack1):

spot on

OpenStudy (anonymous):

Thanks Again!!!

OpenStudy (jack1):

welcome dude